According to data from the US Census Bureau, the population P of Cleveland, Ohio (in thousands) in year x can be approximated b P(x)=.08x^2-13.08x+927, where x=0 corresponds to 1950. In what calendar year, is the population about 1,000,000?
I have tried setting the equation to 1,000,000 (for example: 1,000,000=.08x^2-13.08x+927), but when I solve it, I come out with some outrageous number for the year... could you please show me the steps as to how I am supposed to set it up? I can preform the math, I am just unsure of the set up.

Answer by ankor@dixie-net.com(22740)

According to data from the US Census Bureau, the population P of Cleveland, Ohio (in thousands) in year x can be approximated b P(x)=.08x^2-13.08x+927, where x=0 corresponds to 1950. In what calendar year, is the population about 1,000,000?
I have tried setting the equation to 1,000,000 (for example: 1,000,000=.08x^2-13.08x+927),
:
One fact you overlooked, the population is given "In Thousands" therefore:
1,000,000 would be written as 1000
:
The equation would be:
.08x^2 - 13.08x + 927 = 1000
;
.08x^2 - 13.08x + 927 - 1000 = 0
:
.08x^2 - 13.08x - 73 = 0
Solve this and you will get a positive solution of approx 169 yrs from 1950 which is 2119